The midpoint of AB is (3sqrt5,2sqrt3). If the coordinates of the endpoint B are (-sqrt5,1/2sqrt3), then what are the coordinates of endpoint A?
Answer by stanbon(75887)

The midpoint of AB is (3sqrt5,2sqrt3). If the coordinates of the endpoint B are (-sqrt5,1/2sqrt3), then what are the coordinates of endpoint A?
-----------------------
Let the coordinates of A be (x,y)
--------------------
Note: The coordinates of the midpoint of a line segment are the average
of the coordinates of the endpoints.
---------------------
EQUATIONS:
(x + (-sqrt5))/2 = 3sqrt5
x - sqrt5 = 6sqrt5
x = 7sqrt5
----------------
(y + 2sqrt3)/2 = (1/2)sqrt3
y + 2sqrt3 = sqrt3
y = -sqrt3
-------------------------
The coordinates of A are (7sqrt5 , -sqrt3)
========================
Cheers,
Stan H.